OpenStudy (anonymous):

What is cube root x^3 over 5th root x^2 in simplest form? The right answer is x^3/5. How?

OpenStudy (kamibug):

Is this your problem? \[\frac{ \sqrt[3]{x^3} }{ \sqrt[5]{x^2} }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@KamiBug can you help?

OpenStudy (kamibug):

Let's use an example for the numerator. Say that x = 4 and x^3 = 64 then cube root 64 = 4 So that means cube root x cubed = x For the denominator, just subtract. 5 - 2 = 3 leaves us with cube root x So we have x over cube root x